OpenStudy (anonymous):

passing through (4,4) and parallel to the line whose equation is 2x+y-7=0 slope intercept form

OpenStudy (freckles):

parallel you should think same slope

OpenStudy (freckles):

You need to first find the slope of 2x+y-7=0

OpenStudy (freckles):

this will be the slope of the line you are trying to find.

OpenStudy (freckles):

You are also given a point on this other line that we want to find that point being (4,4)

OpenStudy (freckles):

y-y1=m(x-x1) so you have x1=4 and y1=4

OpenStudy (freckles):

so you only have to find the slope of that line i talked about and replace the m in our other equation

OpenStudy (anonymous):

what happens to -7

OpenStudy (freckles):

put 2x+y-7=0 in y=mx+b form this will give you information to obtain the slope of 2x+y-7=0

OpenStudy (anonymous):

okay but what happens to the -7 or can please take me step by step please

OpenStudy (gorv):

2x+y-7=0

OpenStudy (gorv):

add 7 on both side and tell what u got ?

OpenStudy (anonymous):

2x+y=7

OpenStudy (gorv):

gud

OpenStudy (gorv):

now subtract 2x from both side and tell me what u got ?

OpenStudy (anonymous):

y=7-2x

OpenStudy (gorv):

very good

OpenStudy (gorv):

y=-2x+7

OpenStudy (gorv):

so here it is in standard form

OpenStudy (gorv):

in standard form coefficient of x = slope so slope =??

OpenStudy (anonymous):

2x

OpenStudy (gorv):

coefficeint of x =-2=slope

OpenStudy (gorv):

parallel line will have the same slope

OpenStudy (gorv):

y=-2x+c

OpenStudy (gorv):

now it is passing through (4,4)....so put x=4 and y=4 in above equation and tell methe value of c u got ?

OpenStudy (anonymous):

y=2*(4)+4=12

OpenStudy (gorv):

loll.....y=4 and x=4 u have to put why u put c=4??

OpenStudy (anonymous):

4=8+c

OpenStudy (gorv):

subtract 8 from both side

OpenStudy (anonymous):

-4=c

OpenStudy (gorv):

wait equation was y=-2x+c

OpenStudy (gorv):

4=-2*4+c 4=-8+c

OpenStudy (gorv):

add on each side

OpenStudy (gorv):

c=12

OpenStudy (gorv):

y=-2x+c replace cby its value that will be your line

OpenStudy (anonymous):

okay thanks but how did we get rid of the 7 in the equation?

OpenStudy (freckles):

you don't need the 7

OpenStudy (anonymous):

okay

OpenStudy (freckles):

we know parallel lines have the same slope and different y-intercept

OpenStudy (freckles):

there is nothing we can do with the y-intercept from that equation to determine the y-intercept of the other equation

OpenStudy (freckles):

well except that we know they should be different

OpenStudy (anonymous):

okay thanks for the help